SOHO DAILY REPORT Operations Day - 058/2003 (27 February 03) Mission Day 2645 SOHO FLIGHT OPERATIONS TEAM, NASA/GSFC Ref. #: SH-FOT-DR-030227 ==================================================================== PASS #1 Canberra DSS46 0222-0845 PASS #2 Madrid DSS66 1134-1843 PASS #3 Goldstone DSS27 1829-2130 PASS #4 Goldstone DSS16 2104-059/0208 ********************************************************************* PASS #1 Canberra DSS46 0222-0845 Anomalies: None NRT: None Activities Summary: - Collected MDI-H (0222 - 0237, 0243 - 0825) - OBT Distribution to SWAN and CELIAS - Ranging Data Collection - SSR SEF Reset (x2) - SSR To Record ------------------ PASS #2 Madrid DSS66 1134-1843 Anomalies: G03-0145 1146 3 minutes bad time stamp. TCA problem. DRM101430. G03-0146 1157 6.75 hrs degraded data. Reason unknown. Nonrecoverable. DR#M101431. G03-0147 1808 Wayward command. 7 minute NRT pause. 10 commands retransmitted. NRT: 1509 - 1816 Activities Summary: - SSR Dump - Collected MDI-M (1307 - 1816) - Collected MDI-H (1816 - 1818, 1826 - 1843) - Uplink Delayed Loads: (VIRA30228) - Reaction Wheel Speed Monitoring Daily Adjustment - CRS Drift Estimation Update - Hx limit update - Uplink NOM HGA Table - Ranging Data Collection ------------------ PASS #3 Goldstone DSS27 1829-2130 Anomalies: None NRT: 1907 - 1943 Activities Summary: - Collected MDI-M (1905 - 2100) - Collected MDI-H (1829 - 1905, 2100 - 2130) ------------------ PASS #4 Goldstone DSS16 2104-059/0208 Anomalies: G03-0148 0001 Bad time stamp for 1 hour. TGC problem. DR#G102415. NRT: None Activities Summary: - Collected MDI-M (2150 - 2301) - Collected MDI-H (2104-2140, 2148-2150, 2301-0143, 0149-0208) - Uplink SVM Loads: TTMH_TOMM_MMTOMH_03058 - SSR SEF Reset ------------------ -- End of Report --